# DVLA-RL: Dual-Level Vision-Language Alignment with Reinforcement Learning Gating for Few-Shot Learning

### v1

@@ Abstract
Few-shot learning (FSL) aims to generalize to novel categories with only a few samples. Recent approaches incorporate large language models (LLMs) to enrich visual representations with semantic embeddings derived from class names. However, they overlook progressive and adaptive alignment between vision and language from low-level to high-level semantics, resulting in limited semantic gains. To address these challenges, we propose Dual-level Vision–Language Alignment with Reinforcement Learning gating (DVLA-RL), which consists of Dual-level Semantic Construction (DSC) and RL-gated Attention (RLA). Specifically, DSC conditions LLMs on both class names and support samples to generate discriminative attributes, progressively selects the most relevant ones, and then synthesizes them into coherent class descriptions. This process provides complementary low-level attributes and high-level descriptions, enabling both fine-grained grounding and holistic class understanding. To dynamically integrate dual-level semantics along with the visual network layers, RLA formulates cross-modal fusion as a sequential decision process. A lightweight policy trained with episodic REINFORCE adaptively adjusts the contributions of self-attention and cross-attention to integrate textual and visual tokens. As a result, shallow layers refine local attributes and deep layers emphasize global semantics, enabling more precise cross-modal alignment. This achieves class-specific discrimination and generalized representations with merely a few support samples. DVLA-RL achieves new state-of-the-art performance across nine benchmarks in three diverse FSL scenarios.
@@ 1 Introduction
Deep learning has achieved remarkable success, driven by large-scale labeled datasets(He et al., 2016; Dosovitskiy et al., 2021; Wang et al., 2022; 2023a; Meng et al., 2024; Li et al., 2025a; Sun et al., 2025). However, acquiring such annotations is often costly or even impractical in real-world settings. To reduce the reliance on abundant labeled data and simulate human rapid learning ability from limited experience, Few-Shot Learning (FSL) (Vinyals et al., 2016) is proposed to generalize the knowledge learned from a base set to novel and unseen tasks using a few labeled samples. FSL has garnered significant attention due to its broad application prospects in real-world scenarios, such as rare disease diagnosis and industrial anomaly detection.
@@
To address these limitations, a Dual-level Vision–Language Alignment with Reinforcement Learning gating (DVLA-RL) is proposed to achieve hierarchical and dynamic cross-modal alignment from low-level to high-level semantics. Specifically, DVLA-RL consists of Dual-level Semantic Construction (DSC) and the RL-gated Attention (RLA). In DSC, an LLM is conditioned on both class names and support samples to generate fine-grained attribute candidates. A progressive Top-k selection strategy then iteratively filters these attributes by measuring their semantic relevance to retain only the most discriminative ones, suppressing semantic hallucinations from irrelevant attributes. The selected attributes are further synthesized into coherent class descriptions, providing complementary local and global semantic guidance. To fully integrate these dual-level cues into visual feature extraction, RLA formulates cross-modal fusion as a sequential decision process. A lightweight policy trained with episodic REINFORCE dynamically balances self-attention and cross-attention between visual and textual tokens across network layers. This adaptive mechanism progressively enables shallow layers to focus on fine-grained local details, while deeper layers emphasize holistic contextual semantics. Consequently, DVLA-RL effectively captures complex cross-modal relationships, enhancing both class-specific discrimination and generalization with merely a few support samples.
@@ 2 Related Work
Relying solely on limited visual samples makes it difficult to acquire discriminative features. To mitigate this, semantic-based methods (Xing et al., 2019; Xu and Le, 2022; Chen et al., 2023; Li et al., 2024; Pan et al., 2024; Zhang et al., 2024a; Liu et al., 2025) incorporate additional semantic information from language modality to complement limited visual evidence. Early AM3 (Xing et al., 2019) incorporates class-level semantics into visual prototypes. KTPP (Li et al., 2024) introduces pyramid-shaped cross-modal prompts to guide support features. SIFT (Pan et al., 2024) further employs category-specific semantic embeddings to generate enriched features through an encoding–decoding process. Recent SemFew (Zhang et al., 2024a) and ECER (Liu et al., 2025) explore the use of large language models (LLMs) to construct global abstractions and local structures, but both rely on single-level semantics and static MLP-based fusion modules that lack adaptivity across network depths during visual extraction. In contrast, DVLA-RL achieves hierarchical and dynamic cross-modal alignment by progressively coordinating low-level attributes and high-level descriptions with visual features through an adaptive RL-gated attention mechanism. To the best of our knowledge, this is the first attempt to introduce reinforcement learning for vision–language alignment in few-shot learning.
Figure 1: Overview of the proposed DVLA-RL framework. Dual-level Semantic Construction (DSC) extracts visual attributes with LLMs, progressively selects the most discriminative ones, and synthesizes them into class descriptions. Adaptive RL-Gated Attention (RLA) integrates these dual-level semantics with visual tokens, dynamically balancing self- and cross-attention between visual and textual tokens across layers for hierarchical and adaptive vision–language alignment.
@@ 3.2 Adaptive RL-gated Attention
where $\mathbf{t}^{\star}$ is the 512-dimensional ground-truth text embedding encoded by the CLIP text encoder, and $U$ is a linear projection mapping the fused feature $H$ after GAP to the same text space. The first term promotes visual–text alignment, and the second term measures intra-episode accuracy improvement, with $\mathrm{Acc}_{t}$ and $\mathrm{Acc}_{t-1}$ denoting query accuracy at the current and previous fusion steps. The weights $\lambda_{\text{sim}},\lambda_{\text{imp}}>0$ balance alignment and improvement. The policy gradient is estimated as
@@
where $\pi_{\theta_{\ell}}$ is the policy of the $\ell$-th block. During training, $\alpha$ is applied in a soft–hard mixture (sampling for stop-gradient and averaging for differentiable paths), while at inference it is replaced by the expectation $\mathbb{E}_{\pi_{\theta}}[\alpha]$. Stacking multiple RLA blocks allows the policy to adaptively emphasize attribute-level cues in shallow layers and description-level semantics in deeper layers, enabling hierarchy-aware vision–language alignment in few-shot learning.
@@ 4.3 Model Analysis
GPT-4o Semantics $\bf{82.04}{\scriptstyle\pm.32}$ $87.79{\scriptstyle\pm.28}$ $86.61{\scriptstyle\pm.41}$ $90.07{\scriptstyle\pm.30}$ $\bf{92.45}{\scriptstyle\pm.27}$ $94.27{\scriptstyle\pm.20}$
@@ 5 Conclusion
In this work, we proposed DVLA-RL, a novel framework for hierarchical and adaptive vision–language alignment in Few-Shot Learning (FSL). The Dual-level Semantic Construction (DSC) module progressively extracts fine-grained attributes and synthesizes coherent class descriptions, offering complementary low-level and high-level semantics. To dynamically integrate these semantics, the RL-gated Attention (RLA) module formulates cross-modal fusion as a sequential decision process, adaptively balancing self- and cross-attention between visual and textual tokens with reinforcement learning across network layers. This design enables shallow layers to capture local details and deeper layers to model global context, leading to more discriminative and generalizable representations. Extensive experiments on nine popular benchmarks across three distinct FSL scenarios demonstrate that our DVLA-RL achieves new state-of-the-art results.
@@ Acknowledgment
This work was supported in part by the Natural Science Foundation of China under Grant 62176139 and 62406177, in part by the Shandong Excellent Young Scientists Fund (Oversea) under Grant 2024HWYQ-027, in part by the Natural Science Foundation of Shandong province, China under Grant ZR2023QF124, in part by the Young Scholars Program of Shandong University.
@@ A.3 Comparison with CLIP-based Methods
The CLIP-based FSL line focuses on adapting a powerful vision-language model pretrained on hundreds of millions of image–text pairs, typically through prompt tuning (e.g., CoCoOp [1]) or lightweight adapters (e.g., LDC [2]). In these approaches, the CLIP backbone is fixed or lightly updated, and the performance gains largely stem from leveraging the extensive semantic prior encoded in large-scale CLIP pretraining. Additionally, their tuning scale and procedure closely resemble standard ImageNet-1k classification. In contrast, the proposed method follows the standard episodic FSL paradigm. The backbone is trained from scratch on tiny-scale few-shot datasets such as miniImageNet, where training and test classes are strictly disjoint. As a result, the goals, assumptions, and data regimes of the two FSL settings differ fundamentally.
@@
A.4 More Discussion on Vision–Language Alignment

### v2

@@ Abstract
Few-shot learning (FSL) aims to generalize to novel categories with only a few samples. Recent approaches incorporate large language models (LLMs) to enrich visual representations with semantic embeddings derived from class names. However, they overlook progressive and adaptive alignment between vision and language from low-level to high-level semantics, resulting in limited semantic gains. To address these challenges, we propose Dual-level Vision-Language Alignment with Reinforcement Learning gating (DVLA-RL), which consists of Dual-level Semantic Construction (DSC) and RL-gated Attention (RLA). Specifically, DSC conditions LLMs on both class names and support samples to generate discriminative attributes, progressively selects the most relevant ones, and then synthesizes them into coherent class descriptions. This process provides complementary low-level attributes and high-level descriptions, enabling both fine-grained grounding and holistic class understanding. To dynamically integrate dual-level semantics along with the visual network layers, RLA formulates cross-modal fusion as a sequential decision process. A lightweight policy trained with episodic REINFORCE adaptively adjusts the contributions of self-attention and cross-attention to integrate textual and visual tokens. As a result, shallow layers refine local attributes and deep layers emphasize global semantics, enabling more precise cross-modal alignment. This achieves class-specific discrimination and generalized representations with merely a few support samples. DVLA-RL achieves new state-of-the-art performance across nine benchmarks in three diverse FSL scenarios.
@@ 1 Introduction
Deep learning has achieved remarkable success, driven by large-scale labeled datasets(He et al., 2016; Dosovitskiy et al., 2021; Wang et al., 2022; Wang and others, 2025; Liang et al., 2025; Wang et al., 2023a; Meng et al., 2024; Li et al., 2025a; Wang et al., 2025; Sun et al., 2025; Ma et al., 2025). However, acquiring such annotations is often costly or even impractical in real-world settings. To reduce the reliance on abundant labeled data and simulate human rapid learning ability from limited experience, Few-Shot Learning (FSL) (Vinyals et al., 2016) is proposed to generalize the knowledge learned from a base set to novel and unseen tasks using a few labeled samples. FSL has garnered significant attention due to its broad application prospects in real-world scenarios, such as rare disease diagnosis and industrial anomaly detection.
@@
To address these limitations, a Dual-level Vision-Language Alignment with Reinforcement Learning gating (DVLA-RL) is proposed to achieve hierarchical and dynamic cross-modal alignment from low-level to high-level semantics. Specifically, DVLA-RL consists of Dual-level Semantic Construction (DSC) and the RL-gated Attention (RLA). In DSC, an LLM is conditioned on both class names and support samples to generate fine-grained attribute candidates. A progressive Top-k selection strategy then iteratively filters these attributes by measuring their semantic relevance to retain only the most discriminative ones, suppressing semantic hallucinations from irrelevant attributes. The selected attributes are further synthesized into coherent class descriptions, providing complementary local and global semantic guidance. To fully integrate these dual-level cues into visual feature extraction, RLA formulates cross-modal fusion as a sequential decision process. A lightweight policy trained with episodic REINFORCE dynamically balances self-attention and cross-attention between visual and textual tokens across network layers. This adaptive mechanism progressively enables shallow layers to focus on fine-grained local details, while deeper layers emphasize holistic contextual semantics. Consequently, DVLA-RL effectively captures complex cross-modal relationships, enhancing both class-specific discrimination and generalization with merely a few support samples.
@@ 2 Related Work
Relying solely on limited visual samples makes it difficult to acquire discriminative features. To mitigate this, semantic-based methods (Xing et al., 2019; Xu and Le, 2022; Chen et al., 2023; Li et al., 2024; Pan et al., 2024; Zhang et al., 2024a; Liu et al., 2025) incorporate additional semantic information from language modality to complement limited visual evidence. Early AM3 (Xing et al., 2019) incorporates class-level semantics into visual prototypes. KTPP (Li et al., 2024) introduces pyramid-shaped cross-modal prompts to guide support features. SIFT (Pan et al., 2024) further employs category-specific semantic embeddings to generate enriched features through an encoding-decoding process. Recent SemFew (Zhang et al., 2024a) and ECER (Liu et al., 2025) explore the use of large language models (LLMs) to construct global abstractions and local structures, but both rely on single-level semantics and static MLP-based fusion modules that lack adaptivity across network depths during visual extraction. In contrast, DVLA-RL achieves hierarchical and dynamic cross-modal alignment by progressively coordinating low-level attributes and high-level descriptions with visual features through an adaptive RL-gated attention mechanism. To the best of our knowledge, this is the first attempt to introduce reinforcement learning for vision-language alignment in few-shot learning.
Figure 1: Overview of the proposed DVLA-RL framework. Dual-level Semantic Construction (DSC) extracts visual attributes with LLMs, progressively selects the most discriminative ones, and synthesizes them into class descriptions. Adaptive RL-Gated Attention (RLA) integrates these dual-level semantics with visual tokens, dynamically balancing self- and cross-attention between visual and textual tokens across layers for hierarchical and adaptive vision-language alignment.
@@ 3.2 Adaptive RL-gated Attention
where $\mathbf{t}^{\star}$ is the 512-dimensional ground-truth text embedding encoded by the CLIP text encoder, and $U$ is a linear projection mapping the fused feature $H$ after GAP to the same text space. The first term promotes visual-text alignment, and the second term measures intra-episode accuracy improvement, with $\mathrm{Acc}_{t}$ and $\mathrm{Acc}_{t-1}$ denoting query accuracy at the current and previous fusion steps. The weights $\lambda_{\text{sim}},\lambda_{\text{imp}}>0$ balance alignment and improvement. The policy gradient is estimated as
@@
where $\pi_{\theta_{\ell}}$ is the policy of the $\ell$-th block. During training, $\alpha$ is applied in a soft-hard mixture (sampling for stop-gradient and averaging for differentiable paths), while at inference it is replaced by the expectation $\mathbb{E}_{\pi_{\theta}}[\alpha]$. Stacking multiple RLA blocks allows the policy to adaptively emphasize attribute-level cues in shallow layers and description-level semantics in deeper layers, enabling hierarchy-aware vision-language alignment in few-shot learning.
@@ 4.3 Model Analysis
GPT-4o Semantics $82.04{\scriptstyle\pm.32}$ $87.79{\scriptstyle\pm.28}$ $86.61{\scriptstyle\pm.41}$ $90.07{\scriptstyle\pm.30}$ $\bf{92.45}{\scriptstyle\pm.27}$ $94.27{\scriptstyle\pm.20}$
@@ 5 Conclusion
In this work, we proposed DVLA-RL, a novel framework for hierarchical and adaptive vision-language alignment in Few-Shot Learning (FSL). The Dual-level Semantic Construction (DSC) module progressively extracts fine-grained attributes and synthesizes coherent class descriptions, offering complementary low-level and high-level semantics. To dynamically integrate these semantics, the RL-gated Attention (RLA) module formulates cross-modal fusion as a sequential decision process, adaptively balancing self- and cross-attention between visual and textual tokens with reinforcement learning across network layers. This design enables shallow layers to capture local details and deeper layers to model global context, leading to more discriminative and generalizable representations. Extensive experiments on nine popular benchmarks across three distinct FSL scenarios demonstrate that our DVLA-RL achieves new state-of-the-art results.
@@ Acknowledgment
This work was supported in part by the Natural Science Foundation of China under Grant 62176139 and 62406177, in part by the Jinan Science and Technology Project under Grant 202428008, in part by the Shandong Excellent Young Scientists Fund (Oversea) under Grant 2024HWYQ-027, in part by the Natural Science Foundation of Shandong province, China under Grant ZR2023QF124, in part by the Young Scholars Program of Shandong University.
@@ A.3 Comparison with CLIP-based Methods
The CLIP-based FSL line focuses on adapting a powerful vision-language model pretrained on hundreds of millions of image-text pairs, typically through prompt tuning (e.g., CoCoOp [1]) or lightweight adapters (e.g., LDC [2]). In these approaches, the CLIP backbone is fixed or lightly updated, and the performance gains largely stem from leveraging the extensive semantic prior encoded in large-scale CLIP pretraining. Additionally, their tuning scale and procedure closely resemble standard ImageNet-1k classification. In contrast, the proposed method follows the standard episodic FSL paradigm. The backbone is trained from scratch on tiny-scale few-shot datasets such as miniImageNet, where training and test classes are strictly disjoint. As a result, the goals, assumptions, and data regimes of the two FSL settings differ fundamentally.
@@
A.4 More Discussion on Vision-Language Alignment